### v1

# Optimized Lattice-Structured Flexible EIT Sensor for Tactile Reconstruction and Classification

Abstract
Flexible electrical impedance tomography (EIT) offers a promising alternative to traditional tactile sensing approaches, enabling low-cost, scalable, and deformable sensor designs. Here, we propose an optimized lattice-structured flexible EIT tactile sensor incorporating a hydrogel-based conductive layer, systematically designed through three-dimensional coupling field simulations to optimize structural parameters for enhanced sensitivity and robustness. By tuning the lattice channel width and conductive layer thickness, we achieve significant improvements in tactile reconstruction quality and classification performance. Experimental results demonstrate high-quality tactile reconstruction with correlation coefficients up to 0.9275, peak signal-to-noise ratios reaching 29.0303 dB, and structural similarity indexes up to 0.9660, while maintaining low relative errors down to 0.3798. Furthermore, the optimized sensor accurately classifies 12 distinct tactile stimuli with an accuracy reaching 99.6%. These results highlight the potential of simulation-guided structural optimization for advancing flexible EIT-based tactile sensors toward practical applications in wearable systems, robotics, and human-machine interfaces.
Index Terms:
Electrical impedance tomography (EIT), lattice structure, tactile reconstruction, tactile classification
I Introduction
Electronic skin (e-skin) is an emerging sensing technology that employs flexible and stretchable materials integrated with embedded sensors to replicate the sensory functions of human skin[1]. By enabling robots to interact with and accurately perceive their surroundings, even in complex and unstructured environments [2], e-skin represents a promising pathway toward embodied intelligence. Therefore, it has attracted significant attention for applications in human-machine interfaces (HMI) [3], health monitoring [4], and wearable devices [5]. Typically, e-skin systems consist of a flexible substrate integrated with various sensors capable of detecting stimuli such as pressure, temperature, and humidity. Among these, tactile sensors are particularly crucial for identifying contact locations and quantifying applied forces. Existing tactile sensors are typically based on piezoresistive[6], capacitive[7], optical[8], and magnetic[9] mechanisms. However, these approaches generally utilize arrays of discrete sensing units [10] that require intricate wiring and complex structures, leading to higher production costs, limited flexibility and scalability. Scaling such sensors to larger areas presents challenges such as reduced spatial resolution, higher costs and hardware complexity [11], thereby restricting their practical deployment in real-world applications.
Electrical impedance tomography (EIT) has emerged as a compelling alternative for tactile sensing, addressing many limitations of conventional array-based designs [12]. EIT reconstructs the conductivity distribution within the Region of Interest (ROI) by analysing voltage readouts from sparsely distributed boundary electrodes. This enables the detection of contact locations and applied forces without the need for dense sensor arrays. Its “one-piece” architecture with sparse boundary electrodes offers advantages in terms of scalability, mechanical compliance, and manufacturing simplicity. The integration of flexible materials such as ionic liquids[13], hydrogels[14], and elastic films[15] has further enhanced the stretchability, sensitivity, and self-healing capabilities. Recent developments have demonstrated EIT-based tactile sensors can effectively track dynamic pressure changes[16]. Their capability in capturing temporal variations makes them particularly suitable for applications requiring continuous and responsive tactile sensing.
Conventional EIT-based tactile sensors usually adopt uniformly distributed conductive pathways, resulting in only small potential changes at the boundary electrodes when localised pressure is applied. This inherent characteristic limits the sensitivity of EIT-based tactile sensing systems. Recent research has explored novel structures to improve sensitivity, such as porous[17] and multi-layer structures[18]. However, these approaches often face manufacturing challenges that constrain scalability and durability. Furthermore, the lattice structure facilitates higher current density, enabling more pronounced and rapid conductivity changes under press deformation, thereby significantly improving the responsiveness of the sensing system [19]. However, current lattice-structured EIT tactile sensors face two key limitations[20]. First, critical lattice parameters, such as channel width and conductive layer thickness, are often selected empirically, without systematic optimization. Second, the influence of lattice structure on the EIT signal response has not been quantitatively analyzed or benchmarked against non-lattice configurations. These gaps limit the development of high-sensitivity, robust, and reproducible tactile sensors.
In this work, to resolve the above limitations, we introduce an optimized lattice-structured flexible EIT tactile sensor, where the lattice parameters are systematically refined through three-dimensional coupling field simulations (3D-CFS). This simulation-guided design approach enables a balance of sensitivity, spatial resolution and mechanical robustness. The key contributions of this work are as follows.
  • We develop a flexible lattice-patterned EIT-based tactile sensor fabricated with hydrogel and silicone, whose lattice design parameters (specifically channel width and conductive layer thickness) are optimised through 3D-CFS to maximize sensing performance.
  • We validate the optimized sensor’s capabilities with real-world experiments, demonstrating significantly enhanced sensitivity enabled by the optimized lattice-patterned structure.
  • We demonstrate the optimized sensor’s effectiveness in tactile reconstruction and classification tasks. The sensor achieves superior reconstruction quality and successfully recognizes 12 distinct static and dynamic touch patterns, outperforming comparable EIT-based tactile sensing studies [21, 22, 23].
II Principle of EIT-based tactile sensing
Figure 1: Schematic of lattice-structured flexible EIT sensor for tactile sensing.
Figure 2: Framework of 3D-CFS. (a) 3D EIT tactile sensor modelling. (b) Solid mechanics field. Touch phantom 1 is the example for one touch point, touch phantom 2 is the example for two touch points, and touch phantom 3 is the example for three touch points. (c) Electric current field. E1, …, E15 represent the index of electrodes.
Figure 3: Mean relative voltage changes for 300 random touch phantoms against various lattice parameters.
The principle of EIT-based tactile sensing is illustrated in Fig. 1. When localized pressure is applied to a sample substrate, it induces changes in conductivity within the touched regions, which in turn alter the electric potentials measured at the boundary electrodes. The objective of EIT-reconstruction in tactile sensing is to estimate the pressure-induced conductivity changes $\boldsymbol{\sigma}$ in the sensing region $\boldsymbol{\Omega}$ based on the voltage measurements $\boldsymbol{V}$. EIT involves solving both a forward and an inverse problem [24]. The forward problem predicts surface potentials for a given conductivity and applied current. This relationship is governed by:
$\nabla\cdot\left[{\sigma\left({x,y}\right)\nabla u\left({x,y}\right)}\right]=0,\quad\left({x,y}\right)\in\boldsymbol{\Omega}$ (1)
where $\sigma\left({x,y}\right)$ represents the conductivity at $\left({x,y}\right)$, and $u\left({x,y}\right)$ is the potential distribution. This can be linearized as follows:
$\Delta\boldsymbol{V}=\boldsymbol{J}\Delta\boldsymbol{\sigma}$ (2)
where $\Delta$ denotes discrete change and $\boldsymbol{J}$ is the Jacobian matrix. The EIT inverse problem involves estimating $\Delta\boldsymbol{\sigma}$ from $\Delta\boldsymbol{V}$, which is typically formulated as a regularized optimization problem:
$\arg\min_{\Delta\boldsymbol{\sigma}}\frac{1}{2}\|\boldsymbol{J}\Delta\boldsymbol{\sigma}-\Delta\boldsymbol{V}\|_{2}^{2}+\lambda R(\Delta\boldsymbol{\sigma})$ (3)
where $R$ is the regularization term incorporating prior knowledge, and $\lambda>0$ is the regularization factor that balances data fidelity and regularization.
In contrast, learning-based approaches for solving the EIT inverse problem aim to establish an inverse mapping operator, denoted as $\mathcal{F}$, using data-driven methods:
$\Delta\hat{\sigma}=\mathcal{F}(\Delta V)$ (4)
III Lattice-patterned Flexible Tactile Sensor
III-A 3D-CFS
We conducted 3D-CFS to determine the optimal layer thickness and lattice channel width for maximizing sensor sensitivity. The 3D-CFS framework combines a solid mechanics field to model the deformation caused by external pressure under different touch patterns, with an electric current field model that simulates the injected current through electrode pairs positioned along the model boundaries. The adjacent-drive adjacent-measurement excitation protocol was employed to obtain 104 voltage measurements for each touch pattern.
As shown in Fig. 2(a), the simulated sensor model is designed to be 100 × 100 × 9 mm and consists of three layers and 16 evenly distributed boundary electrodes. The substrate and sealing layers are non-conductive, and the sealing layer has a fixed thickness of 2 mm. The thickness of the substrate layer is changed with the intermediate conductive lattice layer to maintain a constant total model thickness of 9 mm. To evaluate and optimise sensor sensitivity, six different lattice channel widths and five different conductivity layer thicknesses were simulated to assess the influence of lattice channel widths and conductive layer thickness on sensor sensitivity.
As shown in Fig. 2(b), three different touch conditions are applied by varying the number of touch points from one to three. In each condition, 100 sets of touch phantoms are randomly generated, with press depths from 1 mm to 5 mm and touch radius from 6.25 mm to 8.75 mm.
To determine the optimal lattice configuration, the average relative voltage change $V^{\text{rel}}$ was introduced to quantify the sensor’s electrical response to external tactile stimuli.
$V^{\text{rel}}=\frac{1}{N}\sum_{i=1}^{N}\frac{|V_{i}^{\text{touch}}-V_{i}^{\text{ref}}|}{|V_{i}^{\text{ref}}|}$ (5)
where $N$ is the total number of measurement channels; $V_{i}^{\text{touch}}$ represents the voltage measured at the i-th channel under touch conditions, and $V_{i}^{\text{ref}}$ is the corresponding voltage under the reference (no-touch) condition.
By averaging the normalized voltage differences across all channels, this metric captures the overall sensitivity of the sensor configuration to pressure-induced conductivity changes.
Different lattice parameters were calculated, with the conductive layer thickness from 1 mm to 5 mm and the channel width from 0 mm (no lattice pattern) to 6 mm. The result is as shown in Fig. 2d, each average value represents the average relative voltage change from 300 random touch phantoms across three touch conditions. The results demonstrate that a conductive layer thickness of 3 mm and a channel width of 4 mm consistently provided the highest sensitivity.
III-B Sensor Fabrication
Figure 4: Hydrogel fabrication process. (a) The solution contained 21.5 weight percent (wt%) lithium chloride (Sigma-Aldrich), 8 wtacrylamide (Sigma-Aldrich), 32.27 wtdeionised water and 37 wtethylene glycol (Sigma-Aldrich) was prepared. Ethylene glycol maintains hydrogel pliability across a wide temperature range, essential for stable sensor performance. (b) The solution was stirred at 60°C and 700 rpm for two hours to achieve uniform mixing and dissolution. (c) To initiate a polyacrylamide-based network, 1 wtpolyacrylamide (Sigma-Aldrich) was introduced, providing mechanical stability and enhancing sensor durability. (d) The solution was stirred overnight at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete dissolution and formation of a uniform polymer network. (e) A crosslinking agent, 0.15 wtN,N’-methylenebisacrylamide (Sigma-Aldrich), was added to control the mechanical strength of the gel. Additionally, 0.08 wt2-hydroxy-2-methylpropiophenone (Sigma-Aldrich), a photoinitiator, was added to enable UV curing of the hydrogel. (f) The mixture was shielded from light and stirred for an additional two hours at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete homogenization.
Figure 5: Fabrication process of the lattice-patterned tactile sensor. (a) A 3D-printed mould A was prepared to define the sensor structure. (b) A pre-gel mixture of silicone was poured into the mould. This mixture was prepared by combining parts A and B of Ecoflex 00-30 (Smooth-On Inc.) in a 1:1 ratio. (c) After allowing the silicone to cure for 4 hours at room temperature, the substrate layer was released from the mould. (d) A 3D-printed mould B, containing designated regions for the sensing area, cable channels, and electrode holes, was prepared. Cables and electrodes were installed within the mould. (e)The released substrate layer was assembled with the sealed 3D-printed mould B, and the gap was sealed with silicone. (f) The hydrogel precursor was poured into the designated sensing area and cured under UV light (365 nm) for two hours. (g) A final layer of silicone pre-gel was applied over the entire sensor to form a sealing layer, protecting it from environmental interference. (h) The fabricated lattice-patterned tactile sensor.
Based on the optimal design parameters identified through the 3D-CFS simulations, a conductive layer thickness of 3 mm and a lattice channel width of 4 mm, we proceeded to fabricate the lattice-structured EIT tactile sensor accordingly. The sensing layer is fabricated from a hydrogel primarily composed of acrylamide, a hydrophilic monomer known for polyacrylamide-based hydrogels, which retains water and ensures sensor functionality in humid conditions. The inclusion of lithium chloride, a common ionic conductor, further enhances the hydrogel’s ionic conductivity. The hydrogel precursor synthesis procedure is shown in Fig. 4.
As shown in Fig. 5, the hydrogel and silicone layers were integrated to fabricate the complete tactile sensor (Fig. 5h). To enhance durability and maintain long-term functionality, the entire structure was encapsulated with a silicone layer to prevent moisture loss from the hydrogel.
III-C Real-world Sensitivity Characterization and Validation
Figure 6: Sensitivity characterization of different sensors. (a) Mean relative voltage change of three sensors. (b) Tactile reconstruction using Tikhonov Regularization [25] for lattice-patterned and uniform sensors. (c) Sensor cycling test of the lattice-patterned sensor for 100 pressing cycles.
To verify the optimized sensor parameters identified through CFS, we fabricated three sensors using the same procedures (refer to Fig. 5) and a consistent conductive layer thickness of 3 mm: two lattice-patterned sensors with different lattice parameters and one without the lattice structure for comparison. These sensors were subjected to a continuously increasing localized pressure with strains ranging from 0 mm to 4 mm, as illustrated in Fig. 6a. The pressure was applied through a circular indenter with a radius of 8.75 mm, simulating the typical contact area of a human fingertip. The pressing location, marked by the blue dot in the inset of Fig. 6a, remained consistent across all measurements. The results demonstrate that the sensor with optimised lattice parameters (channel width of 4 mm) consistently shows superior sensitivity compared to the non-lattice structured sensor and lattice-structure sensor with non-optimal parameters (channel width of 6 mm), aligning well with the simulation results. Specifically, at the press depth of 4 mm, the mean relative voltage change for the sensor with the optimized lattice (channel width of 4 mm) reaches approximately 0.0692, compared to  0.0613 for the 6 mm lattice sensor and only  0.0375 for the non-lattice sensor. This corresponds to an 84.5% improvement over the non-lattice structure and a 12.9% enhancement compared to the non-optimal parameters (channel width of 6 mm), clearly confirming the effectiveness of the optimized lattice pattern in enhancing sensitivity.
Additionally, tactile reconstruction results comparing the optimised lattice-patterned sensor and the sensor without the lattice channel are illustrated in Fig. 6b. The reconstructed images show that the optimised lattice-patterned sensor successfully reconstructs tactile information even at lower compressive strains, whereas the sensor without the lattice structure fails to identify tactile input at these shallow press depths. Furthermore, as the compressive strain increases, the lattice-patterned sensor consistently demonstrates superior reconstructions compared to the sensor without the lattice structure. These results further validate that the lattice-patterned sensor possesses higher sensitivity and enhanced tactile reconstruction performance.
III-D Cycling Test
To further evaluate the durability and repeatability of the optimized lattice-patterned sensor, we performed a cycling test involving 100 pressing cycles with compressive strains from 0 mm to 4 mm. The contact area and location for the pressure keep the same with Fig. 6a. The results show that the sensor maintains stable output throughout the cycles, with consistent trends in signal changes (Fig. 6c). These results confirm the sensor’s robustness and suitability for applications requiring long-term, repeated tactile interactions.
IV Tactile reconstruction
IV-A Dataset and Network Training
To validate the sensor’s tactile reconstruction performance, we adopted the deep neural network (DNN) architecture in [26]. This model employs the Multilayer Perceptron (MLP) to extract global signal characteristics of voltage measurements, followed by a Convolution Neural Network (CNN) that extracts fine-grained local patterns.
To generate the dataset for the lattice-structured EIT sensor, we constructed a finite element model (FEM) using COMSOL Multiphysics and Matlab. We divide the sensing region with a size of 100 $\times$ 100 mm2 into 2304 elements (48 $\times$ 48) and set the background conductivity $\boldsymbol{\sigma_{0}}$ to 0.00312 S/m. Tactile interactions were simulated using circular regions with diameters ranging from 7.5 mm to 27.5 mm, varying in both location and intensity. The conductivity within these touch regions is randomly assigned values between $0.05\times\sigma_{0}$ and $2\times\sigma_{0}$. The complete dataset comprises 50,000 samples, divided into five subsets containing 10,000 samples each. These subsets correspond to different numbers of touch regions, ranging from 1 to 5 touch areas per sample. Each subgroup is further partitioned into training and validation sets using a 7:3 ratio. To enrich the training dataset by supplementing uncollected location information, we used a data augmentation approach to convert a single EIT measurement into 8 distinct readouts [26]. To further improve generalisation, additive Gaussian noise with Signal-to-Noise Ratios (SNRs) of 35, 40, 45, 50, 55, 60 and 65 dB was added to the seven augmented measurements, respectively. Finally, this process yielded a total of 280,000 samples for training and 15,000 samples for validation.
For DNN training, we retained the parameter settings and optimization strategy reported in [26]. To match the target resolution of 48 $\times$ 48 in this study, an upsampling layer was appended to rescale the original 64 $\times$ 64 output.
IV-B Reconstruction Results and Discussion
Figure 7: Tactile reconstruction using experiment data. Images are normalised.
Figure 8: Tactile reconstruction of unseen real-world complex touches. Images are normalised.
As shown in Fig. 7, from Phantom 1 to Phantom 8, we applied 1 to 5 contacts at different positions using a linear robotic platform (DLE-RG-0001). The reconstruction results demonstrate that the sensor with the DNN model successfully reconstructs various touch patterns with good accuracy. To quantitatively assess reconstruction quality, we evaluate the results using correlation coefficients (CC), peak signal-to-noise ratio (PSNR), relative image error (RE), and structural similarity indexes (SSIM). Across all phantoms, the model achieves strong performance, with CC values reaching up to 0.8856, RE values as low as 0.4028, SSIM values up to 0.9648, and PSNR values reaching 29.0303 dB. These results indicate that the proposed sensor and model combination delivers high-fidelity reconstructions with strong consistency and visual clarity.
Beyond the above benchmark scenarios, we further validated the system’s performance on real-world measurements involving more complex and natural touch patterns. These include multi-finger contact, rectangular contact regions, and annular touch profiles. As shown in Fig. 8, the DNN model effectively reconstructs these diverse tactile configurations. Supplementary Video 1 further demonstrates the model’s real-time tactile reconstruction capabilities on selected phantoms, confirming its suitability for practical, interactive scenarios.
IV-C HMI Application
Figure 9: HMI application: controlling the Super Mario Bros Game.
To demonstrate the practical utility of the proposed tactile sensor, we implemented it as a human–machine interface (HMI) for controlling a virtual character in the Super Mario Bros game, as shown in Fig. 9. The left column shows the physical tactile contact on the sensor, and the right column describes the actions generated in the virtual environment, illustrating the character’s response to each tactile input. When the user presses the sensor in different locations, the corresponding action is performed in the game, such as advancing or jumping. When the user presses the sensor for a different time, it will also be recognised as a distinct motion amplitude, such as low-altitude jumping and high-altitude jumping. Supplementary Video 2 showcases real-time interactions with the Super Mario Bros Game. These results confirm the feasibility of using our tactile sensors for real-time, intuitive control in virtual environments, underscoring their ability to capture and transform subtle tactile inputs for multi-functional HMI applications.
V Tactile Classification
V-A Dataset and Network Training
To validate the superior spatio-temporal characteristics of our sensor, we conducted continuous motion detection experiments using various tactile interactions. Specifically, we selected tactile gestures with different movement trajectories as reference actions to assess the system’s performance. As shown in Fig. 10, the dataset includes 12 distinct interaction types such as no contact, finger press, four-finger scratch, fist press, finger double tap, palm pat, finger swipe up, finger swipe down, finger swipe left, finger swipe right, zoom-in and zoom-out. Each interaction type comprises 1000 samples, each containing 15 frames of voltage measurements. Then, we divided the 12,000 signals into a training set and a validation set in an 8:2 ratio, with 90% of the data used for model training and 10% for model validation and tuning. To fully evaluate the generalisation and robustness, 1,200 new signals were collected to be the test dataset, with each interaction type comprising 100 samples.
To validate the effectiveness and generalizability of the input data, we implemented two types of models: a CNN-based architecture adapted from [21, 22, 23], and a Long Short-Term Memory (LSTM) model. By comparing the performance of both models, we aim to assess the suitability of our data for both spatial and temporal learning approaches. The model parameters are listed in Table I.
TABLE I: Model parameters
Parameter CNN LSTM
Optimizer type Adam Adam
Base learning rate 1.0e-4 1.0e-4
Total epoch 200 200
The epoch of Best Model 60 72
Loss function Cross Entropy Loss Cross Entropy Loss
Batch size 256 256
Input dimension 1$\times$15$\times$104 15$\times$104
Output dimension 12 12
Hardware RTX 4080 GPU RTX 4080 GPU
Activation function ReLU -
Conv layer 1 (Channels, Kernel) [1→32, 5×5] -
MaxPool 1 (Kernel, Stride) 2×2, 2 -
Conv layer 2 (Channels, Kernel) [32→16, 5×5] -
MaxPool 2 (Kernel, Stride) 2×2, 2 -
Flattened size 1248 -
FC layer 1 (Input, Output) [1248, 128] -
FC layer 2 (Input, Output) [128, 12] -
LSTM layer (Input, Hidden, Layers) - [104, 64, 2]
FC layer (Input, Output) - [64,12]
Figure 10: 12 classes of tactile stimulation.
V-B Classification Results and Discussion
Figure 11: Confusion matrices of tactile classification result. (a) The confusion matrix for the LSTM model. (b) The confusion matrix for the CNN model.
As shown in the confusion matrices in Fig. 11, the sensor system effectively distinguishes between 12 tactile modalities based on analysing 1200 samples. The CNN-based model achieves a classification accuracy of 98.5%, while the LSTM-based model further improves the performance, reaching an accuracy of 99.6%. Table II presents a comparative summary with prior studies. While earlier works based on CNN architectures handled fewer touch categories (4 to 10 classes) with accuracies ranging from 93.3% to 98.7%, our method extends the classification to 12 categories without compromising accuracy. These results highlight the robustness and reliability of our approach in handling complex touch interaction scenarios.
TABLE II: Comparison of various studies in classifying touch modalities
Study Number of Classes Model Accuracy
Park et al. [21] 4 CNN 98.7%
Yang et al. [22] 9 93.3%
Yang et al. [23] 10 95.3%
Ours 12 CNN 98.5%
LSTM 99.6%
Conclusion
We developed a lattice-structured flexible EIT tactile sensor, systematically optimized via 3D coupling field simulations and validated through real-world experiments. The sensor demonstrates high sensitivity, spatial resolution, and durability, achieving a CC of 0.9275, PSNR of 29.03 dB, SSIM of 0.9660, and RE as low as 0.3798, in reconstructing various touch patterns. It also accurately classifies 12 tactile gestures with over 99.6% accuracy, confirming its effectiveness for both tactile reconstruction and interpretation. These results highlight its strong potential for HMI and wearable applications. Future work will focus on scaling the design and integrating multi-modal sensing to support next-generation electronic skin systems.